A proposed project has an estimated cost of $3,500 to construct a mall booth and will produce an estimated annual OCF of $450 for 10 years. At the end of 10 years, the mall booth will be removed at an estimated cost of $310. Calculate the proposed project's estimated NPV using a WACC of 10%.
Answer #1

Let the cash flows in year n be denoted by CFn

Given,

CF0 = -3500
CF1 = 450
CF2 = 450
........
CF10 = 450 - 310

WACC = r = 10%

Hence, NPV = -CF0 + CF1/(1+r) + CF2/(1+r)2 ..... CF10/(1+r)10

= -3500 + 450/(1+0.10) + 450/(1+0.10)2 + ...... 450/(1+0.10)10 - 310/(1+0.10)10

= -3500 + 450[1- (1+0.10)-10]/0.10 - 310/(1+0.10)10

= - $ 854.46